QUESTION 1 1. The sample variance or the sample standard deviation are good approximate of the population variance or standard deviation? True or False

QUESTION 2 1. The Central limit theorem states that the individual results or when n is 1 in an experiment that unique outcome follow a Normal Distribution? True or False

QUESTION 3 1. In hypothesis testing alpha is the probability of being judged correct? True or False

QUESTION 4 1. If the critical Z is ± 1.96 and the experimental Z is 1.0 that means that the null hypothesis is rejected? True or False

QUESTION 5 1. In hypothesis testing with a sample size of 25, standard deviation of 2, and the difference between the sample mean and the population mean of +3.0 (positive 3), the Z is positive 7.5? True or False
Answer #1

1.1) True

2.1) False

3.1) False

4.1) False

Since the test statistic value is not greater than the positive critical value (1 < 1.96), so the null hypothesis is not rejected.

5.1) True

z = (エー)/(ulo)

= 3/(2/\sqrt 25)

= 7.5